use the ordered pair shown on the graph (97.6,2.2)(2000,13.5) to find the slope of the line to the nearest hundreath
Found 2 solutions by stanbon, Alan3354:
Answer by stanbon(75887)

use the ordered pair shown on the graph (97.6,2.2)(2000,13.5) to find the slope of the line to the nearest hundreath
-----
slope = (13.5-2.2)/(2000-97.6) = 11.3/1902.4 = 0.0059
---
Round to nearest hundreth you get: 0.006
===========================================
Cheers,
Stan H.
=========

Answer by Alan3354(69443)

use the ordered pair shown on the graph (97.6,2.2)(2000,13.5) to find the slope of the line to the nearest hundreath
==========================
slope = diffy/diffx
m = (13.5-2.2)/(2000 - 97.6)
m = 11.3/1902.4
m = 0.0059
m = 0.01 to the nearest 100th